# Improving Coverage in Combined Prediction Sets with Weighted p-values

### v2

@@ Appendix D A note on computational complexity
The complexity of prediction set aggregation methods was first observed by Cherubin, who compares the computational overhead of their majority vote aggregation method with the overhead of p-value aggregation methods. They note that while majority vote requires simpler operations to determine whether each label is included in the final prediction set, it also requires that predictions be recomputed for each significance level, making it less efficient when sets must be constructed at multiple thresholds. In contrast, p-value methods allow the aggregation to be computed once and then applied to any significance level without additional computation.
@@ Appendix E Further implementation details
E.1 Synthetic dataset
For our ablative/expository experiments, we generate a simple homoskedastic dataset to simulate a regression task. Each input is a 16-dimensional vector drawn from a standard normal distribution, and the output label is the sum of the feature values with additive Gaussian noise. Specifically, for each sample $X_{i}$, we have
@@ E.1 Synthetic dataset
E.2 Mixture-of-experts model
In all of our experiments, we use an MoE model of $N$ experts, where each expert is a linear model defined as $f_{i}(x;\theta_{i})$. Here, $x$ represents the input features and $\theta_{i}$ the parameters of the $i$th expert. Each expert considers a different subset of features depending on the experiment (§E.4 and §B.1.2). The experts are trained independently using L-BFGS to minimize mean square error (MSE) on the training data.
The routing network is also a linear model, and is responsible for generating a set of weights $\{w_{i}(x)\}_{i=1}^{N}$ that determine the contribution of each expert to the final prediction. To ensure that the weights are positive and sum to one, softmax is applied to the routing network outputs:
$w_{i}(x)=\frac{\exp(g_{i}(x;\phi))}{\sum_{j=1}^{N}\exp(g_{j}(x;\phi))},$
where $g_{i}(x;\phi)$ denotes the output of the routing network for the $i$th expert, and $\phi$ represents the parameters of the routing network.
The routing network is trained to minimize MSE of the aggregated prediction
$\ell_{routing}=\frac{1}{|S_{train}|}\sum_{i\in S_{train}}\left(y_{i}-\sum_{j=1}^{N}w_{j}(x_{i})f_{j}(x_{i};\theta_{j})\right)^{2}.$
In our experiments, the trained MoE model is utilized in two ways. For split conformal, we use the weighted sum of experts as a black-box predictive model. For weighted aggregation, we use the weights from the routing network to scale the p-value functions of each expert.
E.3 Computing prediction sets
To compute prediction sets, we leverage what we know about the structure of the p-value function. Observe that for a fixed input $x$, the function $\hat{p}_{k}(x,y)$ (2) is piecewise constant with discontinuities only at the values of $y$ that solve $s_{k}(x,y)=s_{k}(X_{i},Y_{i})$. (For an absolute residual score function $s_{k}(x,y)=|y-\hat{\mu}_{k}(x)|$, these values are equal to $\hat{mu}_{k}(x)\pm s(X_{i},Y_{i})$ for $i\in S_{k}$.) For each test point $x$, $\hat{p}$ is therefore a step function of $y$ with finite discontinuities, and the weighted average p-value function is likewise a step function of $y$, with its discontinuities as a union of the $K$ separate sets of discontinuities.
We use the portion library [39] to represent and manipulate our prediction sets; this also allows us to compute their Lebesgue measures without any discretization.
E.4 Expert feature assignment for UCI experiments
@@
For the UCI experiments, the features for each dataset are partitioned into groups of semantically related features, and each expert in the MoE specializes in a single group of features. We list these groups and their features below, where the feature names are provided with the original dataset [40].

### v3

@@ Appendix D A note on computational complexity
The complexity of prediction set aggregation methods was first observed by cherubin2019majority, who compares the computational overhead of their majority vote aggregation method with the overhead of p-value aggregation methods. They note that while majority vote requires simpler operations to determine whether each label is included in the final prediction set, it also requires that predictions be recomputed for each significance level, making it less efficient when sets must be constructed at multiple thresholds. In contrast, p-value methods allow the aggregation to be computed once and then applied to any significance level without additional computation.
@@ Appendix E Further implementation details
E.1 Data
Synthetic data
For the ablation/expository experiments in §B.1, we generate a simple homoskedastic dataset to simulate a regression task. Each input is a 16-dimensional vector drawn from a standard normal distribution, and the output label is the sum of the feature values with additive Gaussian noise. Specifically, for each sample $X_{i}$, we have
@@ E.1 Data
Data split
In all experiments, we randomly partition the data into four disjoint subsets of train, calibration, merge, and test. For the experiments on UCI data, we allocate 400 samples for train/calibration/merge with a 50/40/10 split (where the number of training samples follows barber2021predictive). For methods that do not require a merge set (e.g. split conformal), we use a 50/50 train/cal split over the same 400 sample budget. This ensures that all methods use the same amount of training data.
E.2 Mixture-of-experts (MoE) model and optimization
Model architecture
To isolate the effects of aggregation from model complexity, we use linear models for both the experts and the routing networks.
We define an MoE model with $K$ experts, where the $i$th expert is a linear model $f_{i}(x;\theta_{i})$ parameterized by $\theta_{i}$, and the routing network is a linear model parameterized by $\phi$. For each expert $i$, it produces a raw output $g_{i}(x;\phi)$. To ensure the expert weights $\{w_{i}(x)\}_{i=1}^{K}$ are positive and sum to one, we apply a softmax function to these outputs:
$w_{i}(x)=\frac{\exp(g_{i}(x;\phi))}{\sum_{j=1}^{K}\exp(g_{j}(x;\phi))},$
These weights determine the contribution of each expert to the final prediction.
Feature assignments
@@ E.2 Mixture-of-experts (MoE) model and optimization
In the UCI experiments, the features for each dataset are partitioned into semantically related groups, and each expert specializes in a single group (see §E.5 for feature groups). In our ablation experiments on synthetic data, we vary the degree of feature overlap between experts to explicitly study its impact on coverage in §B.1.2.
Optimization and sequential training
All models are optimized using L-BFGS to minimize mean squared error (MSE). We adopt a sequential training procedure to improve stability [nie2021evomoe, lin2024moe]. First, we train each expert independently on its assigned subset of features. Then, we train the routing network with the full MoE to minimize MSE of the aggregated prediction
$\ell_{\text{routing}}=\frac{1}{|S_{\text{train}}|}\sum_{i\in S_{\text{train}}}\left(y_{i}-\sum_{j=1}^{K}w_{j}(x_{i})f_{j}(x_{i};\theta_{j})\right)^{2}.$
Note that, while experts may observe different features, all components are trained on the same train set.
Routing at inference
In our experiments, the trained MoE model is utilized in two ways. For split conformal, we use the weighted sum of experts as a standard black-box predictor. For weighted aggregation, we use the weights from the routing network to scale the p-value functions of each expert.
E.3 Implementation of weighted aggregation
Computing prediction sets
We leverage the structure of the p-value function to efficiently compute prediction sets.
For a fixed input $x$, the function $\widehat{p}_{k}(x,y)$ (2) is piecewise constant, with discontinuities occurring only at values of $y$ that solve $s_{k}(x,y)=s_{k}(X_{i},Y_{i})$. (For an absolute residual score function $s_{k}(x,y)=|y-\widehat{\mu}_{k}(x)|$, these values correspond to $\widehat{\mu}_{k}(x)\pm s(X_{i},Y_{i})$ for all $i\in S_{k}$.) Thus, for each test point $x$, $\widehat{p}_{k}$ is a step function of $y$ with finite discontinuities, and the weighted average p-value function is also a step function of $y$, with its discontinuities as a union of the $K$ separate sets of discontinuities. We compute prediction sets by identifying these candidate discontinuities, evaluating the p-value at the midpoint of each interval, then taking the union of all intervals where $\widehat{p}(x,y)>\alpha$.
We use the portion library [portion] to represent prediction sets and efficiently compute their Lebesgue measures without discretization.
Computing correction factors $\widehat{m}^{*}$, $\widehat{m}^{\dagger}$, $\widehat{m}^{\ddagger}$
All correction factors are computed using the empirical CDF of the aggregated p-values on the merging set $S_{\text{merge}}$.
For each point $(X_{i},Y_{i})\in S_{\text{merge}}$, we calculate the aggregate p-value function $P_{\text{all}}^{i}=p_{\text{all}}(X_{i},Y_{i};\mathbf{W}^{(i)})$. We then construct the empirical CDF $\widehat{F}$ and compute the ratios $\widehat{F}(P_{\text{all}}^{i})/P_{\text{all}}^{i}$ for all points in the set.
  • $\widehat{m}^{*}$ is the maximum of these ratios over $i\in S_{\text{merge}}$.
  • $\widehat{m}^{\dagger}$ (targeted) restricts this maximum to points where $\widehat{F}(P_{\text{all}}^{i})\leq\alpha^{\prime}$.
  • $\widehat{m}^{\ddagger}$ (precise) uses only the first point $i$ where $\widehat{F}(P_{\text{all}}^{i})\geq\alpha^{\prime}$.
Formal definitions for $\widehat{m}^{*}$ and $\widehat{m}^{\dagger}$ are provided in §A.5.
Role of the merging set $S_{\text{merge}}$
The merging set is used to construct an empirical CDF in order to estimate the correction factor $m^{*}$ that makes the aggregated p-value valid. The size of this set induces a trade-off between giving a conservative or precise estimate, which we demonstrate experimentally in §B.1.1.
A small $|S_{\text{merge}}|$ results in a noisier empirical CDF; to maintain validity, the correction factor tends to be large, leading to overconservative prediction sets. As $|S_{\text{merge}}|$ increases, the empirical CDF becomes closer to the true CDF, allowing for a more precise estimate of $m^{*}$ and coverage closer to target. Our ablation study in Figure 4 indicates that a modest sample size of $\approx 160$ is sufficient to reduce overcoverage to within $3\%$, while maintaining validity.
E.4 Evaluation metrics
Worst-slice (WS) coverage
We report WS coverage to evaluate the local validity of our prediction sets.
A slab of the feature space is a region bounded by hyperplanes:
$S_{v,a,b}=\{x:a\leq v^{\top}x\leq b\}.$
WS coverage is the minimum empirical conditional coverage over all slabs of the feature space that contain at least some fraction $\delta$ of the test points:
$\min_{v,a,b:\mathbb{P}\{X\in S_{v,a,b}\}\geq\delta}\mathbb{P}\{Y\in C(X)|X\in S_{v,a,b}\}.$
Intuitively, WS coverage measures performance on the most challenging slice that still has a sufficient number of test samples. In our experiments, we approximate this metric by evaluating coverage over 1000 random projections $v$ and slab boundaries $(a,b)$ and reporting the minimum, following romano2020classification.
E.5 Expert feature assignments for UCI experiments
For the UCI experiments, the features for each dataset are partitioned into groups of semantically related features, and each expert in the MoE then specializes in a single group of features. We list these groups and their features below, where the feature names are provided with the original dataset [frank2010uci].